Find the "algebraic solution" of the following:
a) i^45
b) i^425
how do I know if I'm finding the "algebraic solution" or just the solution? I don't know what they mean by "algebraic solution"? I'm taking an Algebra course so how could it not be "algebraic"?

Answer by stanbon(75887)

Find the "algebraic solution" of the following:
a) i^45 = i^[4*11+1] = (i^4)^11*i = 1*i = i
--------------------------------------------------
b) i^425 = i^[4*106+1] = (i^4)^106*i = 1*i = i
----------------
Cheers,
Stan H.